# NViST: In the Wild New View Synthesis from a Single Image with Transformers

### v1

@@ 4.1 MVImgNet
Train/Test Split: MVImgNet contains casually captured videos of over $6.5$ million real-world scenes across $238$ categories. Our training set, contains a subset of $1.14$M frames across $38$K scenes of $177$ categories. For the test set, we hold out every $100^{th}$ scene from each category to a total of $13,228$ frames, from $447$ scenes and $177$ categories.
Pre-processing: MVImgNet utilises COLMAP to estimate camera matrices and generate 3D point clouds. Since each scene has its own scale from SfM, we rescale point clouds to a unit cube, such that the maximum distance along one axis equals $1$.
Additionally, we center the point clouds in the world coordinate system. We downsample the original images by $12$ while preserving the aspect ratio. Camera intrinsics are recalibrated accordingly.
Iterations: We train the model with a single A100 GPU, for over 1 million iterations, which takes approximately one week to complete. During training, the encoder processes about 10 million images, which corresponds to approximately $9$ epochs. For each iteration, we sample $15,000$ rays, and $14,400$ rays are utilised for synthesising the full image to compute the LPIPS loss, while the remaining $600$ rays are employed to sample from different viewpoints.
Implementation Details: The encoder takes images of size $160\times 90$, uses a patch size of $5$, so the total number of feature tokens is $576$. The resolution of the vector-matrix representation is $48$. The patch size for the decoder is $3$, the total number of output tokens $816$, and $16$ heads. The embedding dimension is $768$ for both encoder and decoder, and we sample $64$ points along the ray. The initial learning rates are 2e-5 for encoder, and 1e-4 for decoder and renderer, and we decay them following the half-cycle cosine schedule.
@@
Figure 6: Casual Phone Captures: We demonstrate the performance of NViST in out-of-distribution scenarios (images taken with a commodity phone camera). The first row shows that the model can capture scenes with different categories, and the second row demonstrates how the model works in an outdoor setting. Third and fourth rows show that the model performs well with many objects in the scene.
@@ 4.1.1 Qualitative Results
Results on Test (Unseen) Scenes: As depicted in Figure 4, our model demonstrates its capability to synthesising new views of unknown scenes. Figure 4 highlights the model’s ability to handle objects from diverse categories, such as flowers, bags, helmets and others. We show that NViST can deal with a variety of backgrounds, such as tables, wooden floors, textureless walls or more cluttered environments.
@@
Casual Phone Captures: Our motivation for this paper was to train a feed-forward model that we could easily use on casually captured images, for instance acquired with our own mobile devices. We test the ability of NViST to deal with out-of-distribution scenes/images by performing zero-shot new-view synthesis on scenes captured by us on a mobile-phone device. We pre-process the images in the same manner as MVImgNet, using COLMAP to estimate the focal length and camera distance parameters to be used as conditioning. Figure 6 shows results on out-of-distribution scenes. The top row highlights the model’s ability to process a scene with multiple objects from diverse categories. The second row reveals its competence on outdoor scenes, despite their limited presence in the training set. The third and fourth rows illustrate NViST’s ability to learn scenes with a large number of objects. Figure 1 shows two further examples of results on phone captures.
Depth Estimation: Figure  7 shows qualitative results of the depth predicted by NViST. Since there is no ground-truth depth for these images, we qualitatively compare  NViST with the recent MiDASv3.1 with Swin2-L384 [4]. We provide MiDAS with the GT new view images as input, and scale the predicted depth to have the same range as  NViST. This shows that depth estimation with NViST performs well even though it is not trained with depth supervision, while MiDAS uses direct depth supervision from multiple datasets.
Figure 7: Depth Estimation: We show examples of the depth estimated on test images. Although NViST focuses on novel view synthesis and is trained with RGB losses only, the depth estimation is consistent and finds good object boundaries. We show a comparison with the state-of-the-art disparity estimator MiDAS v3.1 with Swin2-L384 [4]. We provide MiDAS with the GT new view images as input, and scale the predicted depth to the same range as NViST. This shows that depth estimation with NViST performs well even though it is not trained with depth supervision, while MiDAS uses direct depth supervision from multiple datasets.
Figure 8: Qualitative Comparison with PixelNeRF: NViST displays better performance than PixelNeRF [81]. Note that NViST is based on a transformer architecture, uses a perceptual LPIPS loss, and is conditioned on camera parameters to deal with scale ambiguities.
@@
Ours 20.53 0.54 0.35
w/o LPIPS 20.12 0.53 0.52
w/o Camera Conditioning 19.24 0.49 0.41
w/o VM Representation 18.10 0.46 0.50
w/o Updating Encoder 18.24 0.47 0.52
Table 1: Ablation Study: LPIPS loss, conditioning on camera parameters, updating encoder parameters, and VM representation are very effective. For w/o VM Representation, we project features onto three 2D planes and concatenate them.
Comparison with PixelNeRF: We compare with PixelNeRF [81] trained on MVImgNet using the official code released by the authors. Note that we applied the same pre-processing as  NViST. Figure 8 shows that  NViST performs better than PixelNeRF, especially on occluded regions. While PixelNeRF only conditions on spatially aligned features, NViST employs self-attention to deal with occlusions.
4.1.2 Quantitative results
@@
PixelNeRF 17.02 0.41 0.54
Ours 20.53 0.54 0.35
Table 2: Quantitative Evaluations on MVImgNet: We compare NViST with PixelNeRF in the context of novel view synthesis from single images in the wild. NViST performs better than PixelNeRF on all metrics.
Ablation Study: We conducted an ablation study to analyse the effect of design choices on the performance of  NViST as summarised in Table 1. While the perceptual LPIPS [83] loss is not commonly employed in NeRF-based methods, it appears to play a crucial role in improving quality. Conditioning on normalized focal length and camera distance helps the model deal with scale ambiguities. While projecting features onto triplanes has been extensively used before [8, 60, 36], our experiments show that the use of a vector-matrix (VM) representation [12] improves the performance by a huge margin. Updating encoder weights also improves the performance as the output from the encoder is used for cross-attention.
Comparison with PixelNeRF: Table 2 shows a quantitative comparison of NViST and PixelNeRF [81]. NViST outperforms PixelNeRF on every metric, as PixelNeRF works well for nearby views but struggles to deal with occlusions. We could not conduct quantitative comparisons with generative models such as GenVS [9] or with the large-scale model LRM [28] as their models are not publicly available. Moreover, GenVS is category-specific and LRM does not model backgrounds.
@@
Table 3: Quantitative Evaluation on ShapeNet-SRN NViST performs similar to other baselines while we adopt the relative pose. VD stands for Viewset Diffusion. For qualitative comparisons, see supplementary materials.
@@ 4.2 ShapeNet-SRN
ShapeNet-SRN [66] has two categories (cars and chairs) and is a widely used benchmark to compare models that perform novel view synthesis from a single input image. Since all objects are aligned and there is no scale ambiguity, pre-processing is not needed and we do not use the LPIPS loss as it is a synthetic dataset. Table  3 shows that NViST performs similar or better to baseline models on ShapeNet-SRN dataset. Although other models apply absolute pose, we only employ relative pose, which means we do not fully exploit the alignment of objects in ShapeNet-SRN.
@@ 5 Conclusion
In this paper, we introduced NViST, a transformer-based scalable model for new view synthesis from single in-the-wild image inputs. Our evaluations demonstrate robust performance on the MVImgNet test set, novel category synthesis and phone captures containing out-of-distribution scenes. We also validated our design choices and conducted a quantitative comparison on MVImgNet and ShapeNet-SRN. Interesting future directions include applying our model to other vision/graphics tasks such as segmentation or camera/object pose estimation, extending NViST to handle 4D dynamic scenes, scaling up to larger datasets, or integrating it with a latent diffusion model to allow generation or editing of 3D scenes from text.
Limitations: We believe that some of the loss of sharpness in the final renders is due to our computational constraints (only 1 A100 GPU), which led us to downsample images by ×$\times 12$, train on a fraction of the MVImgNet dataset, and reduce the total number of iterations. We pushed for a model with a transformer architecture without adversarial losses or score distillation sampling [51]. This eased and sped up the training, but may have also contributed to some loss of detail in the model outputs.
Acknowledgements: The research presented here has been partly supported by a sponsored research award from Cisco Research. This project made use of time on HPC Tier 2 facilities Baskerville (funded by EPSRC EP/T022221/1 and operated by ARC at the University of Birmingham) and JADE2 (funded by EPSRC EP/T022205/1). We are grateful to Niloy Mitra and Danail Stoyanov for fruitful discussions.
@@ A Implementation Details
Decoder MLPs and Reshaping: The embedding dimension of the decoder is $768$. We have $818$ output tokens, and the channel dimension of VM representation [12] is $32$, with a patch size of $3$ for decoder. For the output tokens corresponding to the matrices $M$ in the VM representation, we deploy MLP to reduce the embedding dimension to $288$. For those corresponding to vectors V, we reduce it to $96$. Subsequently, we reshape them into VM representation.
@@ B Qualitative Results on ShapeNet-SRN
We perform a qualitative comparison with VisionNeRF [38] on ShapeNet-SRN [66] dataset as depicted in Figure  10. VisionNeRF, recognised as one of top-performing models on ShapeNet-SRN, employs ViT [19] as its encoder. Notably, VisionNeRF does not utilise any generative approaches, and was trained using $8$ A100 GPUs. Similarly for MVImgNet, we fine-tune a MAE for the ShapeNet-SRN dataset and initialise the parameters of both encoder and decoder of NViST with this fine-tuned MAE for ShapeNet-SRN. The ShapeNet-SRN images are of resolution $128\times 128$, and we use an encoding patch size of $8$, resulting in $256$ feature tokens. The resolution of VM representation is $64$, and the decoder patch size is $4$, so we use $818$ output tokens, each with an embedding dimension of the Transformer as $768$. We still maintain the relative pose but do not condition on camera parameters as the dataset is aligned and does not have scale ambiguities. We train the model with a single $3090$ GPU with $500,000$ and $700,000$ iterations, respectively for car and chair.

### v2

@@ 4.1 MVImgNet
Train/Test Split: MVImgNet contains videos of over $6.5$ million real-world scenes across $238$ categories. Our training set, contains a subset of $1.14$M frames across $38$K scenes of $177$ categories. For the test set, we hold out every $100^{th}$ scene from each category to a total of $13,228$ frames, from $447$ scenes and $177$ categories.
Pre-processing: MVImgNet uses COLMAP to estimate camera matrices and generate 3D point clouds. Since each scene has its own scale from SfM, we rescale point clouds to a unit cube, such that the maximum distance along one axis equals $1$. Then, we center the point clouds in the world coordinate system. We downsample the original images by ×$\times 12$ while preserving their aspect ratio. Camera intrinsics are recalibrated accordingly.
Implementation Details: NViST has approximately $216$M parameters: $85$M for the encoder, $131$M for the decoder, and $7$K for the renderer. The encoder takes images of size $160\times 90$, using a patch size of $5$, so the total number of feature tokens is $576$. The resolution of the VM representation is $48$. The patch size for the decoder is $3$, the total number of output tokens $816$, and $16$ heads. The embedding dimension is $768$ for both encoder and decoder, and we sample $48$ points along the ray. We train NViST with 2$\times$ A100-40GB GPUs for approximately one week, using a batch size of $22$ images and rendering $330$K pixels, up to one million iterations. The initial learning rates are 6e-5 for encoder, and 4e-4 for decoder and renderer and we decay them following the half-cycle cosine schedule.
@@
Figure 6: Casual Phone Captures: We demonstrate NViST in OOD scenarios. First row: the model can capture different categories. Second row: outdoor setting. Third row: many objects.
@@ 4.1.1 Qualitative Results
Results on Test (Unseen) Scenes: As depicted in Figure 4, our model demonstrates its capability to synthesise new views of unknown scenes. Figure 4 highlights the model’s ability to handle objects from diverse categories, such as flowers, bags, helmets and others. We show that NViST can deal with a variety of backgrounds, such as tables, wooden floors, textureless walls or more cluttered environments.
@@
Casual Phone Captures: Our motivation for this paper was to train a feed-forward model that we could easily use on casually captured images, for instance acquired with our own mobile devices. We test the ability of NViST to deal with out-of-distribution scenes/images by performing zero-shot new-view synthesis on scenes captured by us on a mobile-phone device. We pre-process the images in the same manner as MVImgNet, using COLMAP to estimate the focal length and camera distance parameters to be used as conditioning. Figure 6 shows results on out-of-distribution scenes. The top row highlights the model’s ability to process a scene with multiple objects from diverse categories. The second row reveals its competence on outdoor scenes, despite their limited presence in the training set. The third row illustrates NViST’s ability to learn scenes with a large number of objects. Figure 1 shows two further examples of results on phone captures.
Depth Estimation: Figure  7 shows qualitative results of the depth predicted by NViST. Since there is no ground-truth depth for these images, we qualitatively compare  NViST with the recent MiDASv3.1 with Swin2-L384 [6]. As MiDAS predicts depth from RGB images, we provide the GT novel view image as input. This shows that depth estimation with NViST performs well even though it is not trained with depth supervision, while MiDAS uses direct depth supervision from multiple datasets.
Figure 7: Depth Estimation: Examples of the depth estimates on test images. Although NViST focuses on novel view synthesis and is trained with RGB losses only, depth estimation is consistent and finds good object boundaries. We show a comparison with the state-of-the-art disparity estimator MiDAS v3.1 with Swin2-L384 [6]. We provide MiDAS the GT new view images as input, as it cannot do novel view synthesis. NViST performs well even though it is not trained with depth supervision, unlike MiDAS.
Figure 8: Qualitative Comparison with PixelNeRF and VisionNeRF: NViST displays better performance than PixelNeRF [99] and VisionNeRF [46], especially when the target view is far away from the input view.
@@
PixelNeRF [99] 17.02 0.41 0.54
VisionNeRF [46] 19.82 0.51 0.47
Ours 20.83 0.57 0.29
Table 1: Quantitative Tests on MVImgNet: We compare NViST with PixelNeRF and VisionNeRF on new view synthesis from single in-the-wild images. NViST outperforms both on all metrics.
4.1.2 Comparisons with baseline models
We compare NViST with PixelNeRF [99] and VisionNeRF [46], all trained on MVImgNet using their official code releases and applying the same pre-processing as NViST. Figure 8 and Table 1 show that  NViST outperforms both models by a large margin. Both models rely on aligned features, but have limitations dealing with occlusion. We qualitatively compare with the pre-trained Zero-1-to-3 [49] on a phone capture in Figure  9, using the approximate camera pose as it assumes a centered object. Since Zero-1-to-3 and its follow-up works assume a 3DoF camera and do not model the background, we could not conduct quantitative comparisons.
Figure 9: Qualitative comparison with Zero-1-to-3 [49] on a phone captured input image: Zero-1-to-3 shows result w/masked input, Zero1to3 shows result w/full image input. None result in good novel-view synthesis, except NViST.
We could not conduct quantitative comparisons with generative models such as GenVS [13] or with the large-scale model LRM [33] as their models are not publicly available. Moreover, GenVS is category-specific and LRM does not model backgrounds.
@@ 4.1.2 Comparisons with baseline models
Ours 20.83 0.57 0.29
w/o LPIPS 20.72 0.54 0.46
w/o Camera Conditioning 20.24 0.49 0.36
Concat Camera Parameters(1) 20.81 0.57 0.30
w/ Self-Attention Decoder(2) 20.74 0.57 0.31
w/o VM Representation(3) 19.60 0.49 0.44
w/o Updating Encoder 18.54 0.47 0.49
Table 2: Ablation: For (1) we concatenate high dimensional camera feature tokens to output tokens instead of adaptive layer normalisation. For (2) we update all tokens via self-attn (no cross-attn). For (3), we use triplane instead of VM as the representation.
4.1.3 Ablation Study
We conducted an ablation study to analyse the effect of design choices on the performance of  NViST as summarised in Table 2. While the perceptual LPIPS loss [101] is not commonly employed in NeRF-based methods, it appears to play a crucial role in improving quality. Conditioning on normalized focal length and camera distance helps the model deal with scale ambiguities, and adaptive layer normalisation performs better than concatenating camera parameters to output tokens. Instead of employing cross-attention in Figure 2, we can concatenate feature tokens and output tokens and update both of them using self-attention. However, it increases memory consumption and does not lead to a better result in Table 2. While projecting features onto triplanes has been extensively used before [12, 70, 44, 33], our experiments show that the use of a vector-matrix (VM) representation [16] improves performance. Note that for the triplane representation, we use a 1-layer MLP to regress occupancy, while occupancy is directly calculated using Equation 5 in our VM representation. Updating encoder weights also improves the performance as the encoder output is used for cross-attention.
@@ 4.1.3 Ablation Study
Table 3: Quantitative Evaluation on ShapeNet-SRN: NViST performs similarly to other baselines on ShapeNet Cars/Chairs, although our method only requires relative pose. For qualitative comparisons, see supplementary materials.
@@ 4.2 ShapeNet-SRN
ShapeNet-SRN [79] has two categories (cars and chairs) and is a widely used benchmark to compare models that perform novel view synthesis from a single input image. Since all objects are aligned and there is no scale ambiguity, pre-processing is not needed and we do not use the LPIPS loss as it is a synthetic dataset. Table  3 shows that NViST performs similarly to baseline models on ShapeNet-SRN dataset. Although other models apply absolute pose, we only employ relative pose, which means we do not fully exploit the alignment of objects in ShapeNet-SRN.
@@ 5 Conclusion
We have introduced NViST, a transformer-based scalable model for new view synthesis from a single in-the-wild image. Our evaluations demonstrate robust performance on the MVImgNet test set, novel category synthesis and phone captures of out-of-distribution scenes. Our design choices were validated via ablations and a quantitative comparison was conducted on MVImgNet and ShapeNet-SRN. Interesting future directions include extending NViST to adopt a probabilistic approach and to multiview inputs.
Limitations: Some loss of sharpness could be due to our computational constraints, which led us to downsample images by ×$\times 12$ and train on a fraction of the original dataset. We pushed for a transformer-based architecture, without GAN losses or SDS [60], which eased and sped up training, but may have also contributed to some loss of detail.
Acknowledgements: Our research has been partly supported by a sponsored research award from Cisco Research and has made use of time on HPC Tier 2 facilities Baskerville (funded by EPSRC EP/T022221/1 and operated by ARC at the University of Birmingham) and JADE2 (funded by EPSRC EP/T022205/1). We are grateful to N. Mitra and D. Stoyanov for fruitful discussions.
@@ A Implementation Details
Decoder MLPs and Reshaping: The embedding dimension of the decoder is $768$. We have $818$ output tokens, and the channel dimension of VM representation [16] is $32$, with a patch size of $3$ for the decoder. For the output tokens corresponding to the matrices $M$ in the VM representation, we deploy MLP to reduce the embedding dimension to $288$. For those corresponding to vectors V, we reduce it to $96$. Subsequently, we reshape them into VM representation.
@@ B Qualitative Results on ShapeNet-SRN
We perform a qualitative comparison with VisionNeRF [46] on ShapeNet-SRN [79] dataset as depicted in Figure  12. VisionNeRF, recognised as one of top-performing models on ShapeNet-SRN, employs ViT [24] as its encoder. Notably, VisionNeRF does not utilise any generative approaches, and was trained using $8$ A100 GPUs. Similarly for MVImgNet, we fine-tune a MAE for the ShapeNet-SRN dataset and initialise the parameters of both encoder and decoder of NViST with this fine-tuned MAE for ShapeNet-SRN. The ShapeNet-SRN images are of resolution $128\times 128$, and we use an encoding patch size of $8$, resulting in $256$ feature tokens. The resolution of VM representation is $64$, and the decoder patch size is $4$, so we use $818$ output tokens, each with an embedding dimension of the Transformer as $768$. We still maintain the relative pose but do not condition on camera parameters as the dataset is aligned and does not have scale ambiguities. We train the model with a single $3090$ GPU with $500,000$ and $700,000$ iterations, respectively for car and chair.
Figure 11: Qualitative Results on Test (Unseen) Scenes of MVImgNet [100]: NViST can synthesize high-quality novel view on challenging scenes from single in-the-wild input images.